OpenStudy (anonymous):

Evaluate: \[\lim_{x \rightarrow -\infty} Sin(x) dx\] just to be clear it Does Not Exist right?

OpenStudy (turingtest):

right

OpenStudy (anonymous):

I'd like to introduce you to WolframAlpha for all your computational needs!

OpenStudy (turingtest):

of course, just thinking about the graph of sinx you don't need wolf to solve it

OpenStudy (anonymous):

This is true, just pointing out that straight-forward answer checking can be done via Mr. Alpha.

OpenStudy (anonymous):

sine is periodic, so it takes on every value between -1 and 1 infinitely often, hence no limit

OpenStudy (anonymous):

follow up question: \[\lim_{t \rightarrow -\infty} \int\limits_{t}^{-\pi/2} \cos \theta d \theta\] is divergent corret?

OpenStudy (turingtest):

yep

OpenStudy (anonymous):

mucho gracias

OpenStudy (turingtest):

de nada